OpenStudy (kenkaneki12280000):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y –2 4 2 6 Line 2 x y –2 –2 0 4 Line 1: y =------ x + -------- Line 2: y =-------x+ --------

OpenStudy (kenkaneki12280000):

@flowey15 @Englishguy @Astrophysics @alivejeremy @FaiqRaees @haleyelizabeth2017

OpenStudy (faiqraees):

Can you find the slope for the first equation?

OpenStudy (kenkaneki12280000):

i dont know how ;-;

OpenStudy (kenkaneki12280000):

@MeganXOXO @alivejeremy

OpenStudy (alivejeremy):

The slope is found by y_{2} -y_{1} / x_{2} - x_{1} Line 1: y2 is 6, y1 is 4, x2 is 2, x1 is -2 6-4/2- -2= 2/4=1/2 Slope intercept is written as y= mx+b. m is the slope, and b is the intercept (the intercept is the value of y when x is 0). To find out the intercept, we plug in values of y, m and x 4= 1/2(-2) +b 4=-1+b 5=b The equation for line 1 is y= 1/2x +5 Using the same formulas as above, the equation for line 2 is y= 3x+4 The solution to the system is a point where both lines would meet in the graph, or a set of x and y values that satisfies both equations. y= 3x+4 y= 1/2x+5 3x+4 = 1/2x+5 3x= 1/2x +5 -4 3x= 1/2x +1 3x-1/2x= 1 2.5x=1 2.5x/2.5=1/2.5 x=0.4 Plug 0.4 into the slope intercept equation y= 3(0.4) +4 y= 1.2 +4 y=5.2 y= 1/2(0.4) +5 y= 0.2 +5 y= 5.2 The solution is 0.4, 5.2 The other set of lines Line 1: y= 3x+2 Line 2: y= -2x +7 Solution: 1,5

OpenStudy (kenkaneki12280000):

@alivejeremy u are to smart for your own good lol thx tho

OpenStudy (alivejeremy):

Yw :)

OpenStudy (kenkaneki12280000):

so that what it means Yw at first i thought it meant yeah right or yea whatever XD

OpenStudy (alivejeremy):

lol yw = Your Welcome

OpenStudy (kenkaneki12280000):

can u help with another